Ryan, Dylan, and Jordan got paid $240 for mowing neighborhood lawns. They split the money in the ratio 5:9:10. how much less did ryan make from Dylan
Answer by ikleyn(52933)

.

The condition means that Ryan got 5x, Dylan 9x and Jordan 10x dollars,

where "x" is (unknown now) the common measure of their earnings.


From the condition,


    5x + 9x + 10x = 240  dollars,   or


    24x           = 240

      x           = 240/24 = 10 dollars.


From this point, complete the solution on your own.